Can someone show me how this is done. I am lost
Six pears and three apples cost $3.90. Two pears and five apples cost $3.30. How much does one pear cost?
Write one equation for each situation and solve the system of equations.
Let p = the cost of one pear
Let a = the cost of one apple
a. Write a system of equations that can be used to determine the cost of pears and apples.
b. Determine the cost of one pear and one apple. Use mathematics to explain how you determined your answer.

Answer by checkley71(8403)

6P+3A=3.90
2P+5A=3.30 NOW MULTIPLY THIS EQUATION BY -3 & ADD THEM
-6P-15A=-9.90
6P+3A=3.90
----------------------------------
-12A=-6
-A=-6/-12
A=$.50 PRICE OF EACH APPLE
6P+3*.5=3.90
6P+1.50=3.90
6P=3.90-1.50
6P=2.40
P=240/6
P=$.40 FOR EACH PEAR.
PROOF
2*.40+5*.50=3.30
.80+2.50=3.30
3.30=3.30